which of the following numbers is the smallest? A. 3/7 B.1/3 C.5/6 D.9/14
Answer by Mathtut(3670)

convert to decimal form and you can see pretty quickly
:
3/7=.4286
1/3=.3333
5/6=.8333
9/14=.6429
:
.3333 is the smallest that is 1/3 letter B is the correct choice